OpenStudy (anonymous):

Find the length of the following segment created by these pairs of endpoints: (0, 1), (6, 6) 7.8 9.1 2.5 6

OpenStudy (anonymous):

find the slope of the equation, then use the distance formula :) Do you know what I'm talking about?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

clue

OpenStudy (anonymous):

!!!!!!!!

OpenStudy (anonymous):

so slope = rise/run = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so it equals (6-1)/(6-0) = 5/6

OpenStudy (anonymous):

sorry I screwd up the distance formula it is actually \[\sqrt{(x _{2} - x _{1})^{2} + y _{2}-y _{1})^{2}}\]

OpenStudy (anonymous):

\[\sqrt{(6-0)^{2}+(6-1)^{2}}\] here is what the equation should look like

OpenStudy (anonymous):

\[\sqrt{(6)^{2}+ (5)^{2}}\] then \[\sqrt{36 +25 } \] next \[\sqrt{61}\] finally\[\sqrt{61}= 7.8(rounded)\] done there is your final answer

OpenStudy (mathmale):

@shadowdragon: thanks for your contribution and desire to be of help, but next time, would you please lead the questioner towards finding his or her own solution? Thanks.